The General Electric Company claims that the life of LED light bulbs are has a mean life of 500 hours with the standard deviation of 50 hours. Test the claim of the manufacture if a random sample of 80 LED light tested and found out to be 450 hours only. Use 5% level of significance.​


1
Expert's answer
2022-04-20T14:40:26-0400

"H_0:\mu=500."

"H_a:\mu \ne 500."

Test statistic: "z=\frac{450-500}{\frac{50}{\sqrt{80}}}=-8.94."

P-value: "p=2P(Z<-8.94)<0.0001."

Since the p-value is less than 0.05, reject the null hypothesis.

There is a sufficient evidence that the claim of the manufacture is not correct.